Jay's favoritebook is 182 pages less than three times Mitchel's favorite book. If both books have the same number of pages, how many pages are in Jay's favorite book?
Answer by ankor@dixie-net.com(22740)

Jay's favoritebook is 182 pages less than three times Mitchel's favorite book. If both books have the same number of pages, how many pages are in Jay's favorite book?
:
It says both books have the same number of pages
Both books = x pages
:
"Jay's book is 182 pages less than three times Mitchel's book."
x = 3x - 182
x - 3x = -182
-2x = -182
x = -182/-2
x = +91 pages
:
:
Check:
91 = 3(91) - 182
91 = 273 - 182